Please Help me!!!
Q:what is the equation of a line parallel to 3y-6x=9,and pass through point (4,8)?
Could you please show me in detail too PLEASE:)
Thank you!!!

Answer by Alan3354(69443)

No reason to panic. Do it like this.
--------------
A line and a point example.
Write in standard form the eqation of a line that satisfies the given conditions. Perpendicular to 9x+3y=36, through (1,2)
-----------------
Find the slope of the line. Do that by putting the equation in slope-intercept form, y = mx + b. That means solve for y.
9x+3y = 36
3y= - 9x + 36
y = -3x + 12
The slope, m = -3
------------------
The slope of lines parallel is the same.
The slope of lines perpendicular is the negative inverse, m = +1/3
----------------
Use y = mx + b and the point (1,2) to find b.
2 = (1/3)*1 + b
b = 5/3
The equation is y = (1/3)x + 5/3 (slope-intercept form)
x - 3y = -5 (standard form)
------------------------
For further assistance, or to check your work, email me via the thank you note, or at Moral Loophole@aol.com